On January 1, 2018 John and Jane decide to admit a new partner, Tom, for a 1/6 interest in the firm for $175,000. The bonus method is used to record the admission of the new partner. After admitting the new partner, the partnership agreement is amended as follows:

Each partner receives 10% interest on his beginning capital balance. Each partner receives an annual salary of $20,000. The residual profit of loss is divided in a ratio of 30% to John, 50% to Jane, and 20% to Tom. Record the journal entry to admission of Tom to the partnership.
Answer #1

For 1/6 th interest in the firm = 175000$

So for 1 = (?)

= 175000*6 = 1050000

20 % of 1050000 = 210000

therefore , bonus = 210000-175000= 35000

Cash A/c Dr 175000

bonus A/c Dr 35000

to partner capital 210000 cr